OpenStudy (anonymous):

Steven has 57 coins in his piggy bank. He only puts pennies and nickels inside. If he has $1.85, how many pennies does he have?

OpenStudy (anonymous):

25 28 31 32

OpenStudy (anonymous):

nickels are 5, pennies are 1 if you put say \(x\) as the number of nickels then they give you \(5x\) cents since there are 57 coins altogether if there are \(x\) nickels there are \(57-x\) pennies, which are obviously on cent each total cents is therefore \[5x+(57-x)\] and you know the total is 185 so set \[5x+(57-x)=185\] and solve for \(x\)

OpenStudy (anonymous):

5x+(57-x)=185 I know i am supposed to start in te parenthesis im just not sure what to do first

OpenStudy (anonymous):

your teacher may have wanted you to write a system of equations put \(x\) as the number of nickels, \(y\) as the number of pennies and then you get two equations \[x+y=57\] and \[5x+y=185\] but then you would solve this by solving the first equation for \(y\) toget \[y=57-x\] substitute it in to the second equation to get \[5x+(57-x)=185\] and so you see that you get the same equation can you solve that one?

OpenStudy (anonymous):

do you subtract 5x from both sides?

OpenStudy (anonymous):

no you should take 4 steps to do this remove parentheses \[5x+57-x=185\] now combine like terms what do you get?

OpenStudy (anonymous):

-6x+57=185

OpenStudy (anonymous):

oops what is \(5-1\) ?

OpenStudy (anonymous):

haha whoops 4x+57=185

OpenStudy (anonymous):

yes now two more steps any ideas or you want me to tell you?

OpenStudy (anonymous):

I think I have to subtract 4x from both sides. ?

OpenStudy (anonymous):

nope that cant be right

OpenStudy (anonymous):

ok lets go slow you want to get \(x\) by itself so you want all the \(x\) on one side of the equal sign and the number on the other so what you need to do is to first subtract \(57\) from both sides

OpenStudy (anonymous):

yes, you no not want to subtract \(4x\) from both sides, that just puts a \(-4x\) on the other side no help

OpenStudy (anonymous):

I swear i was gonna say that first but it didnt look right lol. yay :) 4x+57=180 -57 -57 4x=123 /4 4 x=30.75?

OpenStudy (anonymous):

so 31 :) ?

OpenStudy (anonymous):

well the reason it doesn't look right is because the right hand side was \(185\) not \(180\)

OpenStudy (anonymous):

you are asked for the number of nickels and the number of pennies, so your answer should be a whole number, not a decimal

OpenStudy (anonymous):

haha oh geez. 4x+57=185 -57 -57 4x=128 /4 /4 x=32

OpenStudy (anonymous):

now that looks good therefore there are 32 nickels and also 25 pennies

OpenStudy (anonymous):

thankyoU! :)

OpenStudy (anonymous):

yw